OpenStudy (anonymous):

If tan x°= z/10 and cos x°=10/y, what is the value of sin x°?

OpenStudy (cwrw238):

use the identity tan x = sin x / cos x so sin x = cos x * tan x plug in z/10 and 10/y and simplify

OpenStudy (anonymous):

ok im clueless, how could that even happen?

OpenStudy (cwrw238):

sin x = z / 10 * 10/y the 10's cancel out so sin x = z / y

OpenStudy (cwrw238):

another way to do this is to draw a right angled triangle